Oxygen Optimization stock is currently priced at 59.81 dollars per share. The stock is expected to pay annual dividends that are expected to grow by 2.27 percent per year forever starting after the next dividend is paid in 1 year. The expected return on the stock is 14.07 percent per year. What is the dividend expected to be in 8 years?
Answer #1
Step-1:Calculation of growth rate
Required return = (D1/P0)+g
0.1407 = (D1/59.81)+0.0227
0.1180 = D1/59.81
          7.06 = D1
Step-2:Calculation of dividend in 8 years
D8 = D1*(1+g)^n Where,
= 7.06*(1+0.0227)^7 D8 = Dividend in 8 years = ?
= $ 8.26 D1 = Dividend in 1 year = $       7.06
g = Growth rate = 2.27%
n = Time (t8-t1) = 7